It has been found that the average time internet users spend online per week is 18.3 hours.



A random sample of 48 teenagers indicated that their mean amount of internet time per week was 20.9 hours with a population variance of 32.49. At the 0.02 level of significance, can it be concluded that the mean time differs from 18.3 hours per week?

1
Expert's answer
2022-05-10T14:46:46-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=18.3"

"H_1:\mu\not=18.3"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.02," and the critical value for a two-tailed test is "z_c = 2.3263."

The rejection region for this two-tailed test is "R = \{z: |z| > 2.3263\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{20.9-18.3}{5.7/\sqrt{48}}=3.16"

Since it is observed that "|z| = 3.16>2.3263=z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(Z>3.16)= 0.001578,"  and since "p = 0.001578< 0.02=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 18.3, at the "\alpha = 0.02" significance level.